margaret fennell won a prize for her work. she invested part of the money in a certificate of deposit at 4% and $3000 more than that amount in a bond paying 6%. her annual interest income was $780. how much did margaret invest in each rate?
Answer by hamsanash1981@gmail.com(151)

Let the investment amount at 4% interest be x, then amount invested in a bond paying 6% will be x + 3000.
no of years = 1
Simple interest of deposit at 4% = (x * 1* 4)/100 => 4x/100
Simple interest of deposit at 6% = ((x+3000)* 1* 6)/100 => (6x+18000)/100
Total interest = 780
4x/100 + (6x+18000)/100 = 780
4x +6x +18000 = 78000
10x = 60000
x = 6000
therefore amount invested at 4% is 6000 and amount invested at 6% is 9000